### 12 Mar 2024 v1

# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

Abstract
The task of predicting multiple links within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, a challenge increasingly resolvable due to advancements in natural language processing (NLP) and KG embedding techniques. This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms, including chain-of-thought (CoT) prompting and in-context learning (ICL), to enhance multi-hop link prediction in KGs. By converting the KG to a CoT prompt, our framework is designed to discern and learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading Large Language Models (LLMs) within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Our experimental findings discover that integrating ICL and CoT not only augments the performance of our approach but also significantly boosts the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
1 Introduction
Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of three stages. The original KG dataset is loaded in the first stage. In the second stage, we apply the graph preprocessing function to a randomly selected path in the graph, enabling the path to be represented as a chain-of-thought prompt. The third stage involves instruction fine-tuning the model to predict whether the first node is connected to the last node in the sub-KG.
In the realm of data representation and organization, knowledge graphs (KGs) have emerged as a structured and effective methodology, attracting substantial interest in recent years. Despite their utility, gaining valuable insights from KGs, especially when navigating the multi-hop entity relationships, poses a considerable challenge. Although two-node link prediction has yielded promising results, the prediction of multi-hop links remains a difficult task. This difficulty stems from the necessity for models to reason about intermediate entities and their relationships. Further challenge is the issue of debugging KG model predictions, particularly in the context of discriminative prediction, where the model’s lack of explanatory reasoning steps, which obscures the origins of errors, diminishing accuracy and performance. Consequently, the development of models capable of generatively and precisely predicting multi-hop links in KGs is a critical challenge.
Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs [22]. Initially, embedding-based methods were pivotal, utilizing techniques to represent both entities and relations in a KG as low-dimensional vectors to address the link prediction task by preserving the structural and semantic integrity of the graph [2], [24], [5], [8]. As the field progressed, the integration of LLMs began to offer new paradigms, leveraging vast amounts of data and sophisticated architectures to further enhance prediction capabilities and semantic understanding in KGs [1], [32], [33], [31], [17]. This transition shows a significant improvement from using purely mathematical representations to more context-aware methodologies that better understand the knowledge representation.
Despite these successes, our research highlights three major challenges that prior methodologies have not fully addressed, which our approach aims to solve. First, the predominant focus on discriminative, rather than generative, models and outcomes over reasoning processes underscores a gap in the existing methodologies, highlighting the need for models adept at leveraging reasoning to address multi-hop link prediction in KGs. Secondly, existing approaches predominantly focus on predicting links between two immediate nodes, leaving the field of multi-hop link prediction largely unexplored. This limitation affects the models’ ability to navigate and infer connections across extended node sequences. Lastly, the traditional models generally lack emergent abilities, making them less effective when faced with unseen tasks.
To bridge these gaps, we propose the Knowledge Graph Large Language Model Framework (KG-LLM Framework), a novel approach to multi-hop link prediction. As illustrated in Figure 1, envision a KG wherein nodes are interconnected through specific relations. Initially, our framework takes input of the original KG dataset. After preprocessing, a path in the KG (represented by the red links) is transformed into a chain-of-thought (CoT) prompt [27], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of a KG’s CoT prompt, determined by its path length and the number of nodes and connections, directly correlates with the intricacy of the multi-hop problem it represents. Through instruction fine-tuning (IFT) [25] of three Large Language Models (LLMs)—Flan-T5 [25], LlaMa2 [19], and Gemma [18]—our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [29], the model not only improves but also acquires the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
Our study presents the KG-LLM framework as a pioneering approach in the realm of multi-hop link prediction. Our key contributions are:
  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and IFT during training on both ICL and non-ICL evaluations.
  • By reorienting multi-hop link prediction from a traditional classification paradigm to a generative model, our framework significantly simplifies debugging for users.
  • Our findings also indicate that our framework substantially improves the generalizability of LLMs in responding to unseen prompts.
2 Related Work
Recent development of large-scale language models, such as BERT [4], GPT [13], LLaMA [19], Gemini [18] and T5 [14], has significantly advanced NLP in many ways. These models are trained on vast amounts of text data and have demonstrated impressive performance on a wide range of NLP tasks. For example, Flan-T5 (Fine-tuned Language Net Text-To-Text Transfer Transformer) [25], as a variation of T5, is a Transformer-based language model that is pre-trained on a large corpus of diverse text data and fine-tuned on various downstream tasks. The text-to-text training approach makes LLMs particularly suitable for our generative link prediction task, where the model needs to generate the connection between multiple entities based on the input text. Besides, pre-training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [26]. In our study, we will utilize an unseen prompt “What is the relation between Node_id and Node_id?” instead of the original prompt “Is Node_id connected with Node_id?” to evaluate the model’s ability to generalize to unseen prompts.
Another advantage of LLM-based generative modeling is the Chain-of-Thought (CoT) reasoning ability [28], which provides the flexibility of modifying the instruction, options, and exemplars to allow structured generation and prediction. The Chain-of-Thought reasoning process can be naturally integrated with knowledge graphs by translating a reasoning path on a knowledge graph into a natural language style. This flexibility allows us to easily test the model’s ability to follow instructions and make decisions based on the provided information.
Similarly, In-Context Learning (ICL) [3] helps large language models learn from demonstrative examples in the prompt to generate correct answers for the given question. This can also be naturally integrated with knowledge graphs by considering multi-path reasoning in KG. As a result, CoT plus ICL together enable very flexible KG reasoning through natural language, where CoT enables path-based reasoning over KG and ICL further extends the power to multiple paths.
3 Methodology
In this section, we introduce the proposed KG-LLM framework.
3.1 Knowledge Graph Definition
Let $KG=(E,R,L)$ denote a knowledge graph, where $E$ is the set of entities, $R$ is the set of relationships, and $L\subseteq E\times R\times E$ is the set of triples that are edges in the KG. Each triple $(e_{i},r,e_{i+1})\in L$ denotes that there exists a directed edge from entity $e_{i}$ to entity $e_{i+1}$ via the relationship $r$ [23].
3.1.1 Multi-hop Link Prediction
The task of multi-hop link prediction extends beyond simple edge prediction by aiming to identify missing connections over multiple relational steps within a knowledge graph. Specifically, given a sequence of observed triples that form a connected path $P_{obs}={(e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}),\ldots,(e_{n-1},r_{n-1},e_{n})}\subseteq L$, where each triple $(e_{i},r_{i},e_{i+1})$ denotes an observed relation $r_{i}$ between entities $e_{i}$ and $e_{i+1}$. The objective is to predict whether a missing link $l_{miss}=(e_{1},?,e_{n})$ exists by answering True or False [15], [21].
3.1.2 Multi-hop Relation Prediction
The task of multi-hop relation prediction closely aligns the concept of multi-hop link prediction, with a critical distinction in the question and output. Rather than determining the existence of a missing link $l_{miss}$, this task delves into specifying the relationship. This change shifts the focus from a binary existence query to identifying the specific relationship that binds the entities [11].
3.2 Knowledge Prompt
The knowledge prompt refers to a specialized prompt designed for KGs that enables the conversion of a given sequence of observed triples $P_{obs}$ into natural language, thereby facilitating the model learn better. By leveraging the knowledge prompt in the training process, the model is able to more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop link prediction tasks. In this paper, we define the two types of knowledge prompts, standard knowledge prompt and KG-LLM knowledge prompt:
Multi-hop Link Prediction Standard Prompt:
###Input:
[…]
Multi-hop Relation Prediction Standard Prompt:
###Input:
[…]
Multi-hop Link Prediction KG-LLM Prompt:
###Instruction:
[…]
Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
[…]
Multi-hop Relation Prediction KG-LLM Prompt:
###Instruction:
[…]
Answer the question by reasoning step-by-step. Choose from the given options:
[…]
###Input:
[…]
Figure 2: Models processed through the standard framework will be trained using the standard knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks. Standard knowledge prompt stands out for its clarity and simplicity, providing a deep comprehension of the KG and facilitating improved prediction accuracy. On the other hand, KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs. This approach involves breaking down complex and lengthy inputs into manageable, concise processes. By organizing the information systematically, the likelihood of the model misinterpreting the queries is significantly reduced. Such a methodical strategy ensures the model stays highly aware of the Chain-of-Thought (CoT) process throughout its training, thereby supporting its ability to interpret and navigate the KG.
As illustrated in Figure 2, our experiments utilize the previously described prompts: the standard knowledge prompt and the KG-LLM knowledge prompt. The standard knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models. Conversely, the KG-LLM prompt enriches $Q$ with the KG CoT prompt and embeds the CoT reasoning process within $A$. It further introduces an instruction within $Q$ that narrows down the model’s response options (yes/no), thereby guiding it towards more precise outcomes by constraining the range of possible answers.
In addition, we adopt the approach of utilizing one-shot ICL learning, specifically tailored to the FLAN-T5-Large, which is our smallest model. This is because for models of this scale, the impact of utilizing one-shot ICL versus few-shot ICL on accuracy is minimal [3]. To maintain consistency across our experimental framework, we apply the same one-shot ICL methodology to all three models under investigation: FLAN-T5-Large, LlaMa2-7B, and Gemma-7B. This uniform approach ensures that our comparative analysis of the models’ performances is conducted under equivalent learning conditions, thereby providing a fair and controlled assessment of their capabilities.
3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. In the initial stage, the original KG is taken as input. The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM knowledge prompt. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [7] is implemented. This technique strategically adjusts model parameters with minimal computational overhead, thereby optimizing their efficiency and effectiveness in handling complex prediction tasks. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
$\small L=-\sum_{i=1}^{n}\log P(y_{i}|x,y_{1},y_{2},...,y_{i-1})$ (1)
To evaluate our KG-LLM Framework, we train each model twice. As illustrated in Figure 2, the initial training session employs standard knowledge prompt inputs to establish a baseline. Subsequently, we instruction fine-tune the original models using our specialized KG-LLM knowledge prompt inputs. As shown in section 3.2, KG-LLM instruction falls under the classification category. By listing all possible options in the instruction, LLMs can more effectively follow and generate a response from the given choices. For instruction details, please refer to Appendix.
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL), as depicted in Figure 3. The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ emergent ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
Figure 3: This figure illustrates our experimental overflow. We conducted eight tests for each model–Flan-T5, LLaMa2, and Gemma.
4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of the proposed KG-LLM framework to answer the following several key research questions:
$\bullet$ Q1: Which framework demonstrates superior efficacy in multi-hop link prediction tasks in the absence of ICL?
$\bullet$ Q2: Does incorporating ICL enhance model performance on multi-hop link prediction tasks?
$\bullet$ Q3: Is the KG-LLM framework capable of equipping models with the ability to navigate unseen prompts during multi-hop relation prediction inferences?
$\bullet$ Q4: Can the application of ICL bolster the models’ emergent ability in multi-hop relation prediction tasks?
4.1 Experimental Setup
Datasets.
We conduct experiments over $two$ real-world datasets, WN18RR and NELL-995, which are constructed and released by the OpenKE library [6]. WN18RR and NELL-995 are two commonly used benchmark datasets for evaluating knowledge graph embedding models in the field of knowledge representation learning. Statistics of the datasets are shown in Table 1.
Table 1: Basic statistics of the experimental datasets.
Dataset #Entities #Triples # Relations
WN18RR 40,943 86,835 11
NELL-995 75,492 149,678 200
Task splits.
In the preprocessing stage of the WN18RR and NELL-995 datasets, we generate an equal number of positive and negative instances for training. We proceed by randomly extracting 80% of the entries from both the positive and negative datasets to constitute the training data and 20% for testing. To reduce redundancy in the training data, we eliminate duplicates before combining positive and negative training data. During training, we reserve 20% of the training data as validation data. Balancing the ratio of positive and negative data in prediction tasks is based on the premise that it can enhance the overall performance of the model. This approach acknowledges that real-world knowledge graphs often contain a significantly higher number of negative edges (i.e., absence of relationships) compared to positive edges (i.e., presence of relationships). By ensuring equal representation of positive and negative examples, the model is encouraged to learn to predict both types of edges, preventing it from being biased towards only predicting negative edges and improving its performance in predicting positive edges.
Comparing Baselines.
To assess the effectiveness of our KG-LLM framework, we employ models subjected to the Standard framework as baselines and compare their performance across all experiments. In the non-ICL multi-hop link prediction testing, we also compare our LLMs approach with traditional approaches such as TransE [2], Analogy [9], CompleX [20], DistMult [30], and RESCAL [12] to evaluate performance.
Implementation Details.
We trained each model for 2 epochs on an A40 GPU, and despite limited resources, models still showed promising results. In our study, we set the maximum complexity of multi-hop problems to 10 nodes. We also monitor the input token size to optimize processing efficiency, noting that Flan-T5, with its 512-token capacity, had the smallest token size. Consequently, we tailored our experiments to ensure that the maximum length of input data did not exceed 512 tokens.
Table 2: Multi-hop Link Prediction w/o In-Context Learning
Datasets WN18RR NELL-995
Metrics F1 AUC F1 AUC
TransE 0.37 0.37 0.26 0.13
Analogy 0.61 0.52 0.29 0.20
CompleX 0.60 0.51 0.29 0.19
DistMult 0.56 0.48 0.25 0.14
Rescal 0.61 0.50 0.59 0.53
Flan-T5-Large (standard) 0.63 0.67 0.60 0.66
LlaMa2-7B (standard) 0.74 0.72 0.71 0.73
Gemma-7B (standard) 0.76 0.73 0.72 0.71
Flan-T5-Large (KG-LLM) 0.73 0.71 0.70 0.72
LlaMa2-7B (KG-LLM) 0.82 0.83 0.81 0.80
Gemma-7B (KG-LLM) 0.84 0.81 0.82 0.83
Table 3: Multi-hop Link Prediction with In-Context Learning
Datasets WN18RR NELL-995
Metrics F1 AUC F1 AUC
Flan-T5-Large (standard) 0.70 0.74 0.61 0.63
LlaMa2-7B (standard) 0.83 0.85 0.81 0.80
Gemma-7B (standard) 0.86 0.87 0.82 0.83
Flan-T5-Large (KG-LLM) 0.85 0.87 0.78 0.80
LlaMa2-7B (KG-LLM) 0.97 0.95 0.96 0.93
Gemma-7B (KG-LLM) 0.98 0.94 0.95 0.95
Metrics for Multi-hop Link Prediction.
In evaluating the performance of models in multi-hop link prediction tasks, we utilized the Area Under the ROC Curve (AUC) metric [10] and the F1 score [16]. AUC measures the area under the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate against the false positive rate at varying classification thresholds. The threshold is set at a 50% true positive rate and 50% false positive rate, as the number of positive and negative data points are equal in the testing case. A higher AUC value indicates a better ability of the model to differentiate between positive and negative examples. Similarly, the F1 score, ranging from 0 to 1, measures the balance between precision and recall, where higher values represent better performance. For the performance tables presented below, the best performance is indicated in bold, while the second-best performance is indicated with underline.
Metrics for Multi-hop Relation Prediction.
We use accuracy as the performance metric for the multi-hop relation prediction task, which provides an overall measure of the model’s correctness, calculated as the percentage of test cases where the true relation is predicted correctly.
4.2 Multi-hop Link Prediction without In-Context Learning
This section delves into the comparative analysis of the traditional approaches, standard framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios. Table 2 illustrates the baseline performance compared to the results achieved by our innovative approach. Traditional approaches are shown in the top section of the table, LLMs baseline in the middle section, and our framework in the bottom section.
Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and baseline standard framework across both evaluated datasets. This improved performance can be attributed to the KG-LLM framework’s utilization of knowledge prompts. These prompts enable LLMs to take advantage of the intricate network of relationships between entities and their interconnections within the KG. Furthermore, the process of Instruction Fine-Tuning (IFT) precisely tailors the models, honing their focus on the limited options. The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to significantly advance the domain of multi-hop link prediction tasks in real-world applications.
It should be noted that the performance metrics of the LlaMa2 and Gemma models within the standard framework exhibit good performance. This observation suggests that even absent the enhancements introduced by the KG-LLM framework, models can still perform multi-hop link predictions. Nevertheless, the pivotal challenge is overcoming complex multi-hop KG problems. The KG-LLM framework emerges as a decisive factor in this area.
4.3 Multi-hop Link Prediction with In-Context Learning
To evaluate the influence of In-Context Learning (ICL) on models subjected to both the standard and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability. We conducted a comparative analysis using the same LLMs and testing inputs as in the previous section. The key distinction in this evaluation was the incorporation of an ICL example prior to each original testing input. This addition was meticulously designed to ensure task fairness and unbiased assessment; the ICL example shown in the appendix, derived from the training dataset, were restricted to a complexity level of three. This constraint aimed to prevent providing additional knowledge through the ICL example while furnishing a contextually relevant example. By adhering to these principles, we ensured a robust and fair evaluation of model performance under ICL conditions.
Table 3 reveals a notable enhancement in the performance of models under the standard framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%. Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift; notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
An interesting observation emerged regarding the Flan-T5 model under the standard framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration. This phenomenon could be attributed to the increased length and complexity of the testing prompts induced by the added ICL context. While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the standard framework Flan-T5.
Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the standard framework. However, for models refined through the KG-LLM framework, the integration of ICL consistently facilitates performance improvements. This finding underscores the nuanced role of ICL in model optimization, highlighting how effective it is in improving model performance within KG-LLM framework.
Table 4: Multi-hop Relation Prediction w/o In-Context Learning
Datasets WN18RR NELL-995
Metric Accuracy Accuracy
Flan-T5-Large (standard) 0.14 0.06
LlaMa2-7B (standard) 0.23 0.17
Gemma-7B (standard) 0.22 0.19
Flan-T5-Large (KG-LLM) 0.29 0.21
LlaMa2-7B (KG-LLM) 0.33 0.25
Gemma-7B (KG-LLM) 0.36 0.27
Table 5: Multi-hop Relation Prediction with In-Context Learning
Datasets WN18RR NELL-995
Metric Accuracy Accuracy
Flan-T5-Large (standard) 0.35 0.24
LlaMa2-7B (standard) 0.56 0.48
Gemma-7B (standard) 0.58 0.52
Flan-T5-Large (KG-LLM) 0.55 0.53
LlaMa2-7B (KG-LLM) 0.74 0.70
Gemma-7B (KG-LLM) 0.76 0.71
4.4 Multi-hop Relation Prediction without In-Context Learning
In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks emergent ability. We maintain consistency with the testing dataset used in the multi-hop link prediction task to ensure comparability and fairness. The difference lies in the modifications made to the instruction and prompt question presented to the model. Detailed specifications of these instructions are provided in the Appendix.
An intriguing observation emerged from the evaluation, shown in Table 4: all frameworks showcased limited performance in this task. Notably, the KG-LLM framework exhibited marginally superior performance relative to its standard counterpart. Upon examining the predictive outcomes, we reveal a predominant inclination towards negative relation predictions over positive ones, where models generated a ‘no relation’ output for negative instances. This trend was consistent across models operating under both frameworks. The difference in performance between the two frameworks mainly came from the fact that the KG-LLM models were better at accurately identifying some positive relationships.
Answer to Q3: The findings suggest that the KG-LLM framework marginally enhances the models’ emergent abilities. However, it would be premature to assert that our framework equips models with the ability to navigate unseen prompts. This could be attributed to the complexity and difficulty of the new instructions and options. With options no longer limited to a binary yes or no answer, the model may struggle to comprehend the updated instruction and effectively utilize the provided options.
4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the emergent abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.
Answer to Q4: The integration of ICL has markedly improved the models’ ability to engage and excel in unseen tasks, as evidenced by the substantial performance gains detailed in Table 5. The KG-LLM framework, in particular, exhibits a heightened ability to learn and utilize the contextual example provided by ICL, thus significantly improving the accuracy in predicting unseen relations.
5 Conclusions and Future Work
Our study introduces the Knowledge Graph Large Language Model Framework (KG-LLM Framework) as a promising solution for multi-hop generative prediction tasks in knowledge graph analysis. By leveraging techniques such as Chain of Thought (CoT) and Instruction Fine-tuning (IFT), models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks. Overall, the standard framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance. However, the KG-LLM framework models have demonstrated strong performance in multi-hop link prediction tasks, both with and without the utilization of ICL. Additionally, it exhibits a moderate level of ability when aided by ICL in relation prediction tasks that are not seen during model pre-training. These findings highlight the importance of incorporating CoT and IFT methodologies in enhancing the performance of models, while also emphasizing the need for further exploration and improvement in relation prediction tasks and the ability of handling unseen tasks in general.
In our future work, we will consider accessing the model’s reasoning process in the evaluation stage. We are committed to further enhancing KG-related prediction tasks. One key aspect we will focus on is refining the instruction process by limiting the option size, which will help the model better understand and interpret the provided instructions. Additionally, we plan to explore the utilization of prompt quality filters to effectively filter out noisy options, improving the overall accuracy and reliability of the model’s predictions. Furthermore, we will dedicate efforts to advancing our model’s ability to handle tasks that are unseen during pre-training, aiming to enhance its capacity to generalize and handle unseen prompts more effectively. Through these ongoing improvements, we aim to advance the capabilities of KG-LLM models and contribute to the progress of knowledge graph analysis.
Appendix
5.1 Limitations
One limitation of the KG-LLM framework is that we have not tested the model’s performance when the training input exceeds 512 tokens. We anticipate that tasks will become considerably more challenging when dealing with datasets containing numerous relations. Since these relations must be incorporated as options in the instructions, the sheer volume of options can become extensive and potentially mislead the model. To overcome this limitation, future research endeavors to investigate techniques for effectively reducing the size of options. This endeavor will further enhance the efficacy of our KG-LLM framework approach.
5.2 Instruction for Multi-hop Link Prediction
### Instruction:
Below is the detail of a knowledge graph path. Is node [first node id] connnected with node [last node id]?
Answer the question by reasoning step-by-step.
Choose from the given options:
1. Yes
2. No
5.3 Instruction for Multi-hop Relation Prediction
### Instruction:
Below is the detail of a knowledge graph path. What is the relationship between the node [first node id] and the node [last node id]?
Answer the question by reasoning step-by-step.
Choose from the given options:
1. [relation 1]
2. [relation 2]
3. [relation 3]
5.4 Dataset 1
5.4.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
5.4.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
node_47405 has relation_179 with node_46497, means node_47405 music_artist_genre node_46497. node_46497 has relation_180 with node_46501, means node_46497 music_genres_such_as_music_genres node_46501. So node 47405 music_artist_genre node 46501. The answer is yes.
5.4.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. What is the relationship between node_47405 and node_46501?
Answer:
The relationship between the first node and the last node is relation_179.
5.4.4 KG-LLM Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. What is the relationship between node_47405 and node_46501?
Answer:
node_47405 has relation_179 with node_46497, means node_47405 music_artist_genre node_46497. node_46497 has relation_180 with node_46501, means node_46497 music_genres_such_as_music_genres node_46501. So node 47405 music_artist_genre node 46501. The relationship between the first node and the last node is music_artist_genre.
5.5 Dataset 2
5.5.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
The answer is yes.
5.5.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The answer is yes.
5.5.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
The relationship between the first node and the last node is relation_1.
5.5.4 KG-LLM Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The relationship between the first node and the last node is _derivationally_related_form.
5.6 Examples
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_60981 has relation_181 with node_9219. node_9219 has relation_146 with node_31785. node_31785 not connected with node_17868. node_17868 has relation_136 with node_17869. node_17869 not connected with node_23650. node_23650 has relation_11 with node_2407. Is node 60981 connected with node 2407?
Answer from Standard FLAN-T5-Large Model
The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Input:
node_20326 has relation_12 with node_20328. node_20328 not connected with node_8777. node_8777 has relation_71 with node_2906. node_2906 not connected with node_22546. node_22546 has relation_143 with node_10897. node_10897 has relation_2 with node_20. Is node 20326 connected with node 20?
Answer from Standard FLAN-T5-Large Model
The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829. What is the relationship between node_57359 and node_35829 ?
Answer from Standard FLAN-T5-Large Model
The relationship between the first node and the last node is relation_160.
Example of Multi-hop Relation Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. What is the relationship between node_47405 and node_46501 ?
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_58780 has relation_89 with node_5996. node_5996 has relation_1 with node_3. node_3 has relation_11 with node_17475. What is the relationship between node_58780 and node_17475 ?
Answer from Standard FLAN-T5-Large Model
The relationship between the first node and the last node is no.
Table 6: Standard FLAN-T5-Large Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 57359 connnected with node 35829? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_57359 has relation_160 with node_57261, means node_57359 plantincludeplant node_57261. node_57261 has relation_161 with node_35831, means node_57261 plantgrowinginplant node_35831. node_35831 has relation_160 with node_35829, means node_35831 plantincludeplant node_35829. So node 57359 plantincludeplant node 35829. The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… So node 47405 musicartistgenre node 46501. The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path …
1. Yes
2. No
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_57359 has relation_160 with node_57261, means node_57359 plantincludeplant node_57261. node_57261 has relation_161 with node_35831, means node_57261 plantgrowinginplant node_35831. node_35831 has relation_160 with node_35829, means node_35831 plantincludeplant node_35829. So node 57359 plantincludeplant node 35829. The answer is yes.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. What is the relationship between node_9595 and node_6033 ? Answer the question by reasoning step-by-step. Choose from the given options:
1. motherofperson
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_9595 has relation_117 with node_10011, means node_9595 producesproduct node_10011. node_10011 not connected with node_7650 means there is no relationship between node_10011 and node_7650. node_7650 has relation_91 with node_6033, means node_7650 athleteplayssport node_6033. So there is no connection between node 9595 and node 6033. The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path …
1. motherofperson
###Input:
node_47478 has relation_10 with node_1912. node_1912 not connected with node_48232. node_48232 has relation_139 with node_46613.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_47478 has relation_10 with node_1912, means node_47478 latitudelongitude node_1912. node_1912 not connected with node_48232 means there is no relationship between node_1912 and node_48232. node_48232 has relation_139 with node_46613, means node_48232 musicianinmusicartist node_46613. So there is no connection between node 47478 and node 46613. The relationship between the first node and the last node is no.
Table 7: KG-LLM Framework FLAN-T5-Large Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_21118 has relation_2 with node_5. node_5 has relation_1 with node_332. node_332 has relation_73 with node_2405. node_2405 has relation_194 with node_31362. node_31362 has relation_153 with node_22206. node_22206 has relation_154 with node_27734. node_27734 has relation_2 with node_10. Is node 21118 connnected with node 10?
Answer from Standard LlaMa2-7B Model
The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Instruction:
###Input:
node_21567 has relation_141 with node_21540. node_21540 has relation_141 with node_21378. node_21378 has relation_141 with node_21482. node_21482 has relation_141 with node_21391. node_21391 has relation_141 with node_21449. Is node 21567 connnected with node 21449?
Answer from Standard LlaMa2-7B Model
The answer is yes.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722. What is the relationship between node_58493 and node_24722 ?
Answer from Standard LlaMa2-7B Model
The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. What is the relationship between node_47405 and node_46501 ?
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_64967 has relation_37 with node_20088. node_20088 has relation_30 with node_55250. node_55250 has relation_115 with node_11327. What is the relationship between node_64967 and node_11327 ?
Answer from Standard LlaMa2-7B Model
The relationship between the first node and the last node is subpartoforganization.
Table 8: Standard LlaMa2-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 21567 connnected with node 21449? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_21567 has relation_141 with node_21540. node_21540 has relation_141 with node_21378. node_21378 has relation_141 with node_21482. node_21482 has relation_141 with node_21391. node_21391 has relation_141 with node_21449.
Answer from KG-LLM Framework LlaMa2-7B Model
node_21567 has relation_141 with node_21540, means node_21567 chemicalistypeofchemical node_21540. node_21540 has relation_141 with node_21378, means node_21540 chemicalistypeofchemical node_21378. node_21378 has relation_141 with node_21482, means node_21378 chemicalistypeofchemical node_21482. node_21482 has relation_141 with node_21391, means node_21482 chemicalistypeofchemical node_21391. node_21391 has relation_141 with node_21449, means node_21391 chemicalistypeofchemical node_21449. So node 21567 chemicalistypeofchemical node 21449. The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path….
1. Yes
2. No
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033.
Answer from KG-LLM Framework LlaMa2-7B Model
node_9595 has relation_117 with node_10011, means node_9595 producesproduct node_10011. node_10011 not connected with node_7650 means there is no relationship between node_10011 and node_7650. node_7650 has relation_91 with node_6033, means node_7650 athleteplayssport node_6033. So there is no connection between node 9595 and node 6033. The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_20539 has relation_6 with node_10925. node_10925 has relation_120 with node_10922. node_10922 has relation_42 with node_10924.
Answer from KG-LLM Framework LlaMa2-7B Model
node_20539 has relation_6 with node_10925, means node_20539 synonymfor node_10925. node_10925 has relation_120 with node_10922, means node_10925 companyalsoknownas node_10922. node_10922 has relation_42 with node_10924, means node_10922 organizationterminatedperson node_10924. So node 20539 mutualproxyfor node 10924. The relationship between the first node and the last node is relation_7.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_40422 has relation_170 with node_59706. node_59706 has relation_170 with node_59850. node_59850 has relation_170 with node_59582.
Answer from KG-LLM Framework LlaMa2-7B Model
node_40422 has relation_170 with node_59706, means node_40422 professionistypeofprofession node_59706. node_59706 has relation_170 with node_59850, means node_59706 professionistypeofprofession node_59850. node_59850 has relation_170 with node_59582, means node_59850 professionistypeofprofession node_59582. So there is no connection between node 40422 and node 59582. The relationship between the first node and the last node is no.
Table 9: KG-LLM Framework LlaMa2-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033. Is node 9595 connected with node 6033?
Answer from Standard Gemma-7B Model
The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Input:
node_25714 has relation_1 with node_9667. node_9667 has relation_195 with node_2421. node_2421 has relation_73 with node_26004. node_26004 not connected with node_66961. node_66961 has relation_196 with node_24014. node_24014 has relation_1 with node_2156. node_2156 has relation_7 with node_25251. node_25251 not connected with node_4610. node_4610 has relation_56 with node_4806. node_4806 has relation_65 with node_2255. Is node 25714 connected with node 2255?
Answer from Standard Gemma-7B Model
The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_27123 has relation_97 with node_6069. node_6069 has relation_165 with node_6211. node_6211 has relation_165 with node_6163. What is the relationship between node_27123 and node_6163 ?
Answer from Standard Gemma-7B Model
The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. What is the relationship between node_47405 and node_46501 ?
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_49244 has relation_50 with node_1967. node_1967 has relation_50 with node_35902. What is the relationship between node_49244 and node_35902 ?
Answer from Standard Gemma-7B Model
The relationship between the first node and the last node is no.
Table 10: Standard Gemma-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 58493 connected with node 24722? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722.
Answer from KG-LLM Framework Gemma-7B Model
node_58493 has relation_85 with node_2420, means node_58493 politicianrepresentslocation node_2420. node_2420 has relation_7 with node_24721, means node_2420 mutualproxyfor node_24721. node_24721 has relation_145 with node_24722, means node_24721 cityliesonriver node_24722. So there is no connection between node 58493 and node 24722. The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path…
1. Yes
2. No
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722.
Answer from KG-LLM Framework Gemma-7B Model
node_58493 has relation_85 with node_2420, means node_58493 politicianrepresentslocation node_2420. node_2420 has relation_7 with node_24721, means node_2420 mutualproxyfor node_24721. node_24721 has relation_145 with node_24722, means node_24721 cityliesonriver node_24722. So there is no connection between node 58493 and node 24722. The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_1946 has relation_1 with node_3. node_3 has relation_7 with node_346. node_346 has relation_1 with node_2572.
Answer from KG-LLM Framework Gemma-7B Model
node_1946 has relation_1 with node_3, means node_1946 proxyfor node_3. node_3 has relation_7 with node_346, means node_3 mutualproxyfor node_346. node_346 has relation_1 with node_2572, means node_346 proxyfor node_2572. So node 2572 atlocation node 1946. The relationship between the first node and the last node is relation_11.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_42306 has relation_2 with node_21. node_21 has relation_1 with node_332.
Answer from KG-LLM Framework Gemma-7B Model
node_42306 has relation_2 with node_21, means node_42306 atdate node_21. node_21 has relation_1 with node_332, means node_21 proxyfor node_332. So node 332 locationlocatedwithinlocation node 42306. The relationship between the first node and the last node is locationlocatedwithinlocation.
Table 11: KG-LLM Framework Gemma-7B Model Responses to Modified Inputs

### 13 Mar 2024 v2

@@
 1 Introduction
-Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of three stages.
-The original KG dataset is loaded in the first stage.
-In the second stage, we apply the graph preprocessing function to a randomly selected path in the graph, enabling the path to be represented as a chain-of-thought prompt.
-The third stage involves instruction fine-tuning the model to predict whether the first node is connected to the last node in the sub-KG.
 In the realm of data representation and organization, knowledge graphs (KGs) have emerged as a structured and effective methodology, attracting substantial interest in recent years.
@@
 Consequently, the development of models capable of generatively and precisely predicting multi-hop links in KGs is a critical challenge.
+Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of two stages.
+In the first stage, we apply the graph preprocessing function to a randomly selected path in the knowledge graph, enabling the path to be represented as a chain-of-thought prompt.
+The second stage use instruction finetuning to finetune the LLM to predict whether the first node is connected to the last node in the sub-KG.
 Historically, approaches to solving tasks related to KGs can trace their origins from embedding-based methods to more recent advancements with LLMs [22].
@@
 This limitation affects the models’ ability to navigate and infer connections across extended node sequences.
-Lastly, the traditional models generally lack emergent abilities, making them less effective when faced with unseen tasks.
+Lastly, the traditional models generally lack generalization abilities, making them less effective when faced with unseen tasks.
 To bridge these gaps, we propose the Knowledge Graph Large Language Model Framework (KG-LLM Framework), a novel approach to multi-hop link prediction.
@@
 Initially, our framework takes input of the original KG dataset.
-After preprocessing, a path in the KG (represented by the red links) is transformed into a chain-of-thought (CoT) prompt [27], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}.
+After preprocessing, a path in the KG is transformed into a chain-of-thought (CoT) prompt [27], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}.
 The complexity of a KG’s CoT prompt, determined by its path length and the number of nodes and connections, directly correlates with the intricacy of the multi-hop problem it represents.
@@
 As a result, CoT plus ICL together enable very flexible KG reasoning through natural language, where CoT enables path-based reasoning over KG and ICL further extends the power to multiple paths.
+Figure 2: Models processed through the standard framework will be trained using the standard knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
 3 Methodology
@@
 […]
-Figure 2: Models processed through the standard framework will be trained using the standard knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
 The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks.
@@
 For instruction details, please refer to Appendix.
-Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL), as depicted in Figure 3.
+Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL).
 The primary set of inference tasks is centered around multi-hop link prediction.
-Conversely, the secondary set probes the models’ emergent ability in multi-hop relation prediction, particularly with previously unseen prompts.
+Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts.
 Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
-Figure 3: This figure illustrates our experimental overflow. We conducted eight tests for each model–Flan-T5, LLaMa2, and Gemma.
 4 Experiments
@@
 $\bullet$ Q3: Is the KG-LLM framework capable of equipping models with the ability to navigate unseen prompts during multi-hop relation prediction inferences?
-$\bullet$ Q4: Can the application of ICL bolster the models’ emergent ability in multi-hop relation prediction tasks?
+$\bullet$ Q4: Can the application of ICL bolster the models’ generalization ability in multi-hop relation prediction tasks?
 4.1 Experimental Setup
@@
 Consequently, we tailored our experiments to ensure that the maximum length of input data did not exceed 512 tokens.
+Metrics for Multi-hop Link Prediction.
+In evaluating the performance of models in multi-hop link prediction tasks, we utilized the Area Under the ROC Curve (AUC) metric [10] and the F1 score [16].
+AUC measures the area under the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate against the false positive rate at varying classification thresholds.
+The threshold is set at a 50% true positive rate and 50% false positive rate, as the number of positive and negative data points are equal in the testing case.
+A higher AUC value indicates a better ability of the model to differentiate between positive and negative examples.
+Similarly, the F1 score, ranging from 0 to 1, measures the balance between precision and recall, where higher values represent better performance.
+For the performance tables presented below, the best performance is indicated in bold, while the second-best performance is indicated with underline.
+Metrics for Multi-hop Relation Prediction.
+We use accuracy as the performance metric for the multi-hop relation prediction task, which provides an overall measure of the model’s correctness, calculated as the percentage of test cases where the true relation is predicted correctly.
 Table 2: Multi-hop Link Prediction w/o In-Context Learning
@@
 Gemma-7B (KG-LLM) 0.98 0.94 0.95 0.95
-Metrics for Multi-hop Link Prediction.
-In evaluating the performance of models in multi-hop link prediction tasks, we utilized the Area Under the ROC Curve (AUC) metric [10] and the F1 score [16].
-AUC measures the area under the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate against the false positive rate at varying classification thresholds.
-The threshold is set at a 50% true positive rate and 50% false positive rate, as the number of positive and negative data points are equal in the testing case.
-A higher AUC value indicates a better ability of the model to differentiate between positive and negative examples.
-Similarly, the F1 score, ranging from 0 to 1, measures the balance between precision and recall, where higher values represent better performance.
-For the performance tables presented below, the best performance is indicated in bold, while the second-best performance is indicated with underline.
-Metrics for Multi-hop Relation Prediction.
-We use accuracy as the performance metric for the multi-hop relation prediction task, which provides an overall measure of the model’s correctness, calculated as the percentage of test cases where the true relation is predicted correctly.
 4.2 Multi-hop Link Prediction without In-Context Learning
@@
 4.4 Multi-hop Relation Prediction without In-Context Learning
-In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks emergent ability.
+In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks generalization ability.
 We maintain consistency with the testing dataset used in the multi-hop link prediction task to ensure comparability and fairness.
@@
 The difference in performance between the two frameworks mainly came from the fact that the KG-LLM models were better at accurately identifying some positive relationships.
-Answer to Q3: The findings suggest that the KG-LLM framework marginally enhances the models’ emergent abilities.
+Answer to Q3: The findings suggest that the KG-LLM framework marginally enhances the models’ generalization abilities.
 However, it would be premature to assert that our framework equips models with the ability to navigate unseen prompts.
@@
 The ICL example is shown in the appendix.
-The comparative results, as illustrated in Table 5, reveal a significant improvement in the emergent abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4.
+The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4.
 In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.

### 18 Mar 2024 v3

@@
 The task of predicting multiple links within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, a challenge increasingly resolvable due to advancements in natural language processing (NLP) and KG embedding techniques.
-This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms, including chain-of-thought (CoT) prompting and in-context learning (ICL), to enhance multi-hop link prediction in KGs.
-By converting the KG to a CoT prompt, our framework is designed to discern and learn the latent representations of entities and their interrelations.
-To show the efficacy of the KG-LLM Framework, we fine-tune three leading Large Language Models (LLMs) within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation.
+This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms.
+We use our method to convert structured knowledge graph data into natural language, and then use these natural language prompts to finetune large language models (LLMs) to enhance multi-hop link prediction in KGs.
+By converting the KG to natural language prompts, our framework is designed to discern and learn the latent representations of entities and their interrelations.
+To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation.
 Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts.
-Our experimental findings discover that integrating ICL and CoT not only augments the performance of our approach but also significantly boosts the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
+Our experimental findings discover that the performance of our approach can significantly boost the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
 1 Introduction
@@
 Our study presents the KG-LLM framework as a pioneering approach in the realm of multi-hop link prediction. Our key contributions are:
-  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and IFT during training on both ICL and non-ICL evaluations.
-  • By reorienting multi-hop link prediction from a traditional classification paradigm to a generative model, our framework significantly simplifies debugging for users.
+  • By converting knowledge graphs into CoT prompts, our framework allows LLMs to better understand and learn the latent representations of entities and their relationships within the knowledge graph.
+  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and instruction fine-tuning during training.
   • Our findings also indicate that our framework substantially improves the generalizability of LLMs in responding to unseen prompts.
 2 Related Work
-Recent development of large-scale language models, such as BERT [4], GPT [13], LLaMA [19], Gemini [18] and T5 [14], has significantly advanced NLP in many ways.
+Recent development of large-scale language models, such as BERT [4], GPT [14], LLaMA [21], Gemini [20], T5 [15], and Flan-T5 [28] has significantly advanced NLP in many ways.
 These models are trained on vast amounts of text data and have demonstrated impressive performance on a wide range of NLP tasks.
-For example, Flan-T5 (Fine-tuned Language Net Text-To-Text Transfer Transformer) [25], as a variation of T5, is a Transformer-based language model that is pre-trained on a large corpus of diverse text data and fine-tuned on various downstream tasks.
-The text-to-text training approach makes LLMs particularly suitable for our generative link prediction task, where the model needs to generate the connection between multiple entities based on the input text.
-Besides, pre-training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [26].
+The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task, where the model needs to generate the connection between multiple entities based on the input text.
+Recent studies, concurrent work such as GraphEdit [6], MuseGraph [18] and InstructGraph [24], have shown that natural language is effective for representing structural data for LLMs.
+Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [29].
 In our study, we will utilize an unseen prompt “What is the relation between Node_id and Node_id?” instead of the original prompt “Is Node_id connected with Node_id?” to evaluate the model’s ability to generalize to unseen prompts.
@@
 As a result, CoT plus ICL together enable very flexible KG reasoning through natural language, where CoT enables path-based reasoning over KG and ICL further extends the power to multiple paths.
-Figure 2: Models processed through the standard framework will be trained using the standard knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
+Figure 2: Models processed through the ablation framework will be trained using the KG-LLM ablation knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
 3 Methodology
@@
 By leveraging the knowledge prompt in the training process, the model is able to more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop link prediction tasks.
-In this paper, we define the two types of knowledge prompts, standard knowledge prompt and KG-LLM knowledge prompt:
-Multi-hop Link Prediction Standard Prompt:
-###Input:
-[…]
-Multi-hop Relation Prediction Standard Prompt:
-###Input:
-[…]
-Multi-hop Link Prediction KG-LLM Prompt:
+In this paper, we define the two types of knowledge prompts, KG-LLM knowledge prompt and KG-LLM (ablation) knowledge prompt:
+Multi-hop Link Prediction KG-LLM Knowledge Prompt:
 ###Instruction:
@@
 […]
-Multi-hop Relation Prediction KG-LLM Prompt:
+Multi-hop Relation Prediction KG-LLM Knowledge Prompt:
 ###Instruction:
@@
 […]
+Multi-hop Link Prediction KG-LLM(ablation) Knowledge Prompt:
+###Input:
+[…]
+Multi-hop Relation Prediction KG-LLM(ablation) Knowledge Prompt:
+###Input:
+[…]
 The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks.
-Standard knowledge prompt stands out for its clarity and simplicity, providing a deep comprehension of the KG and facilitating improved prediction accuracy.
-On the other hand, KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs.
+KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs.
 This approach involves breaking down complex and lengthy inputs into manageable, concise processes.
@@
 Such a methodical strategy ensures the model stays highly aware of the Chain-of-Thought (CoT) process throughout its training, thereby supporting its ability to interpret and navigate the KG.
-As illustrated in Figure 2, our experiments utilize the previously described prompts: the standard knowledge prompt and the KG-LLM knowledge prompt.
-The standard knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models.
+On the other hand, we remove the instructional component in the KG-LLM (ablation) knowledge prompt and the CoT reasoning process from the expected output.
+This approach stands out for its clarity and simplicity, providing a good comprehension of the KG and facilitating improved prediction accuracy.
+As illustrated in Figure 2, our experiments utilize the previously described prompts: the KG-LLM knowledge prompt and the KG-LLM (ablation) knowledge prompt.
+The KG-LLM (ablation) knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models.
 Conversely, the KG-LLM prompt enriches $Q$ with the KG CoT prompt and embeds the CoT reasoning process within $A$.
@@
 In the initial stage, the original KG is taken as input.
-The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM knowledge prompt.
+The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM and KG-LLM (ablation) knowledge prompt.
 During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B.
@@
 $\small L=-\sum_{i=1}^{n}\log P(y_{i}|x,y_{1},y_{2},...,y_{i-1})$ (1)
-To evaluate our KG-LLM Framework, we train each model twice.
-As illustrated in Figure 2, the initial training session employs standard knowledge prompt inputs to establish a baseline.
-Subsequently, we instruction fine-tune the original models using our specialized KG-LLM knowledge prompt inputs.
-As shown in section 3.2, KG-LLM instruction falls under the classification category.
+To evaluate our KG-LLM Frameworks, we train each model twice.
+As illustrated in Figure 2, the initial training session employs KG-LLM (ablation) knowledge prompt inputs to establish a baseline.
+Subsequently, we instruction fine-tune the original models using KG-LLM knowledge prompt inputs.
+The KG-LLM instruction falls under the classification category.
 By listing all possible options in the instruction, LLMs can more effectively follow and generate a response from the given choices.
@@
 Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts.
-Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
+Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the KG-LLM (ablation) and KG-LLM frameworks.
 4 Experiments
-In this section, we conduct experiments to evaluate the effectiveness of the proposed KG-LLM framework to answer the following several key research questions:
+In this section, we conduct experiments to evaluate the effectiveness of the proposed KG-LLM frameworks to answer the following several key research questions:
 $\bullet$ Q1: Which framework demonstrates superior efficacy in multi-hop link prediction tasks in the absence of ICL?
@@
 Comparing Baselines.
-To assess the effectiveness of our KG-LLM framework, we employ models subjected to the Standard framework as baselines and compare their performance across all experiments.
+To assess the effectiveness of our KG-LLM frameworks, we compare their performance across all experiments.
 In the non-ICL multi-hop link prediction testing, we also compare our LLMs approach with traditional approaches such as TransE [2], Analogy [10], CompleX [22], DistMult [33], and RESCAL [13] to evaluate performance.
@@
 Rescal 0.61 0.50 0.59 0.53
-Flan-T5-Large (standard) 0.63 0.67 0.60 0.66
-LlaMa2-7B (standard) 0.74 0.72 0.71 0.73
-Gemma-7B (standard) 0.76 0.73 0.72 0.71
+Flan-T5-Large (ablation) 0.63 0.67 0.60 0.66
+LlaMa2-7B (ablation) 0.74 0.72 0.71 0.73
+Gemma-7B (ablation) 0.76 0.73 0.72 0.71
 Flan-T5-Large (KG-LLM) 0.73 0.71 0.70 0.72
@@
 Metrics F1 AUC F1 AUC
-Flan-T5-Large (standard) 0.70 0.74 0.61 0.63
-LlaMa2-7B (standard) 0.83 0.85 0.81 0.80
-Gemma-7B (standard) 0.86 0.87 0.82 0.83
+Flan-T5-Large (ablation) 0.70 0.74 0.61 0.63
+LlaMa2-7B (ablation) 0.83 0.85 0.81 0.80
+Gemma-7B (ablation) 0.86 0.87 0.82 0.83
 Flan-T5-Large (KG-LLM) 0.85 0.87 0.78 0.80
@@
 4.2 Multi-hop Link Prediction without In-Context Learning
-This section delves into the comparative analysis of the traditional approaches, standard framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios.
+This section delves into the comparative analysis of the traditional approaches, ablation framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios.
 Table 2 illustrates the baseline performance compared to the results achieved by our innovative approach.
-Traditional approaches are shown in the top section of the table, LLMs baseline in the middle section, and our framework in the bottom section.
-Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and baseline standard framework across both evaluated datasets.
+Traditional approaches are shown in the top section of the table, ablation framework is in the middle section, and KG-LLM framework is in the bottom section.
+Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and ablation framework across both evaluated datasets.
 This improved performance can be attributed to the KG-LLM framework’s utilization of knowledge prompts.
@@
 The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to significantly advance the domain of multi-hop link prediction tasks in real-world applications.
-It should be noted that the performance metrics of the LlaMa2 and Gemma models within the standard framework exhibit good performance.
+We can observe that the performance of traditional models is not ideal.
+This performance is likely attributable to the complexities associated with multi-hop link prediction.
+Unlike the straightforward task of predicting a direct link between two nodes, traditional methods now must consider all intermediate nodes to arrive at a conclusion.
+This additional complexity consequently diminishes their performance.
+It should be noted that the performance metrics of the LlaMa2 and Gemma models within the ablation framework exhibit good performance.
 This observation suggests that even absent the enhancements introduced by the KG-LLM framework, models can still perform multi-hop link predictions.
@@
 4.3 Multi-hop Link Prediction with In-Context Learning
-To evaluate the influence of In-Context Learning (ICL) on models subjected to both the standard and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability.
+To evaluate the influence of In-Context Learning (ICL) on models subjected to both the ablation and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability.
 We conducted a comparative analysis using the same LLMs and testing inputs as in the previous section.
@@
 By adhering to these principles, we ensured a robust and fair evaluation of model performance under ICL conditions.
-Table 3 reveals a notable enhancement in the performance of models under the standard framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%.
+Table 3 reveals a notable enhancement in the performance of models under the ablation framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%.
 Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift; notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
-An interesting observation emerged regarding the Flan-T5 model under the standard framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration.
+An interesting observation emerged regarding the Flan-T5 model under the ablation framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration.
 This phenomenon could be attributed to the increased length and complexity of the testing prompts induced by the added ICL context.
-While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the standard framework Flan-T5.
-Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the standard framework.
+While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the ablation framework Flan-T5.
+Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the ablation framework.
 However, for models refined through the KG-LLM framework, the integration of ICL consistently facilitates performance improvements.
@@
 Metric Accuracy Accuracy
-Flan-T5-Large (standard) 0.14 0.06
-LlaMa2-7B (standard) 0.23 0.17
-Gemma-7B (standard) 0.22 0.19
+Flan-T5-Large (ablation) 0.14 0.06
+LlaMa2-7B (ablation) 0.23 0.17
+Gemma-7B (ablation) 0.22 0.19
 Flan-T5-Large (KG-LLM) 0.29 0.21
@@
 Metric Accuracy Accuracy
-Flan-T5-Large (standard) 0.35 0.24
-LlaMa2-7B (standard) 0.56 0.48
-Gemma-7B (standard) 0.58 0.52
+Flan-T5-Large (ablation) 0.35 0.24
+LlaMa2-7B (ablation) 0.56 0.48
+Gemma-7B (ablation) 0.58 0.52
 Flan-T5-Large (KG-LLM) 0.55 0.53
@@
 An intriguing observation emerged from the evaluation, shown in Table 4: all frameworks showcased limited performance in this task.
-Notably, the KG-LLM framework exhibited marginally superior performance relative to its standard counterpart.
+Notably, the KG-LLM framework exhibited marginally superior performance relative to its ablation counterpart.
 Upon examining the predictive outcomes, we reveal a predominant inclination towards negative relation predictions over positive ones, where models generated a ‘no relation’ output for negative instances.
@@
 The ICL example is shown in the appendix.
-The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4.
+The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the ablation framework and the KG-LLM framework, in contrast to the findings presented in Table 4.
 In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.
@@
 By leveraging techniques such as Chain of Thought (CoT) and Instruction Fine-tuning (IFT), models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks.
-Overall, the standard framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance.
+Overall, the ablation framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance.
 However, the KG-LLM framework models have demonstrated strong performance in multi-hop link prediction tasks, both with and without the utilization of ICL.
@@
 5.4 Dataset 1
-5.4.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
+5.4.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
 ### Context:
@@
 The answer is yes.
-5.4.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
+5.4.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
 ### Context:
@@
 5.5 Dataset 2
-5.5.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
+5.5.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
 ### Context:
@@
 The answer is yes.
-5.5.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
+5.5.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
 ### Context:
@@
 Is node 60981 connected with node 2407?
-Answer from Standard FLAN-T5-Large Model
+Answer from ablation FLAN-T5-Large Model
 The answer is no.
@@
 Is node 20326 connected with node 20?
-Answer from Standard FLAN-T5-Large Model
+Answer from ablation FLAN-T5-Large Model
 The answer is no.
@@
 node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829. What is the relationship between node_57359 and node_35829 ?
-Answer from Standard FLAN-T5-Large Model
+Answer from ablation FLAN-T5-Large Model
 The relationship between the first node and the last node is relation_160.
@@
 node_58780 has relation_89 with node_5996. node_5996 has relation_1 with node_3. node_3 has relation_11 with node_17475. What is the relationship between node_58780 and node_17475 ?
-Answer from Standard FLAN-T5-Large Model
+Answer from ablation FLAN-T5-Large Model
 The relationship between the first node and the last node is no.
-Table 6: Standard FLAN-T5-Large Model Responses to Modified Inputs
+Table 6: ablation FLAN-T5-Large Model Responses to Modified Inputs
 Test Input Models’ Responses
@@
 Is node 21118 connnected with node 10?
-Answer from Standard LlaMa2-7B Model
+Answer from ablation LlaMa2-7B Model
 The answer is yes.
@@
 Is node 21567 connnected with node 21449?
-Answer from Standard LlaMa2-7B Model
+Answer from ablation LlaMa2-7B Model
 The answer is yes.
@@
 node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722. What is the relationship between node_58493 and node_24722 ?
-Answer from Standard LlaMa2-7B Model
+Answer from ablation LlaMa2-7B Model
 The relationship between the first node and the last node is no.
@@
 node_64967 has relation_37 with node_20088. node_20088 has relation_30 with node_55250. node_55250 has relation_115 with node_11327. What is the relationship between node_64967 and node_11327 ?
-Answer from Standard LlaMa2-7B Model
+Answer from Ablation LlaMa2-7B Model
 The relationship between the first node and the last node is subpartoforganization.
-Table 8: Standard LlaMa2-7B Model Responses to Modified Inputs
+Table 8: Ablation LlaMa2-7B Model Responses to Modified Inputs
 Test Input Models’ Responses
@@
 node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033. Is node 9595 connected with node 6033?
-Answer from Standard Gemma-7B Model
+Answer from Ablation Gemma-7B Model
 The answer is no.
@@
 Is node 25714 connected with node 2255?
-Answer from Standard Gemma-7B Model
+Answer from Ablation Gemma-7B Model
 The answer is no.
@@
 node_27123 has relation_97 with node_6069. node_6069 has relation_165 with node_6211. node_6211 has relation_165 with node_6163. What is the relationship between node_27123 and node_6163 ?
-Answer from Standard Gemma-7B Model
+Answer from Ablation Gemma-7B Model
 The relationship between the first node and the last node is no.
@@
 node_49244 has relation_50 with node_1967. node_1967 has relation_50 with node_35902. What is the relationship between node_49244 and node_35902 ?
-Answer from Standard Gemma-7B Model
+Answer from Ablation Gemma-7B Model
 The relationship between the first node and the last node is no.
-Table 10: Standard Gemma-7B Model Responses to Modified Inputs
+Table 10: Ablation Gemma-7B Model Responses to Modified Inputs
 Test Input Models’ Responses